a/1-2a +b/1-2b+c/1-2c=1/2then value of 1/1-2a+ 1/1-2b+1/1-2c=?
 

Grade:11

2 Answers

Arun
5 years ago
Dear student
( – 1/2) [{1 – 1/(1-2a)} + {1 – 1/(1-2b)} + {1 – 1/(1-2c)}] = ½
{1 – 1/(1-2a)} + {1 – 1/(1-2a)} + {1 – 1/(1-2a)} =  1 + 3 = 4
Hope it helps
Shaquib johar
5 years ago
Plz make this math with explaination.